This was a big debate on Facebook and so many people have different answers. Can you please answer this. Thank you!
6/2(2+1) Everyone put either 9 or 1.

Answer by josgarithmetic(39617)

Unsure of the exact rules myself, some people will find the expression unclear.
Do inside parentheses first: 6/2*3

Which way next?
6/2 = 3, then 3*3 =9.
Another way?
Multiplication before division as you move from left to right,
6/(2*3)=6/6=1.

Many people do not like ambiguity of not always remember this order of addition among multiplication and division, and so will clarify what they mean using grouping symbols.